quadratic equation completing the square
x^2+4x-32=0
please work this one for me I'm stumped....answer is 4,-8

Answer by scott8148(6628)

completing the square means making a "perfect square" out of the quadratic ... you can't change the x terms, but the numbers are fair game

x^2+4x=32 ... to make a perfect square; take half of the x coefficient, square it, and add

(4/2)^2=4 ... x^2+4x+4=32+4 ... (x+2)^2=36 ... take square root x+2=+-6 ... so x=4 and x=-8